# KOALA: Empirical Lessons Toward Memory-Efficient and Fast Diffusion Models for Text-to-Image Synthesis

### v2

@@ 1 Introduction
Since Stable Diffusion XL [31] (SDXL) has become the de facto standard model for text-to-image (T2I) synthesis due to its ability to generate high-resolution images and its open-source nature, many models for specific downstream tasks [7; 47; 3; 57; 55] now leverage SDXL as their backbone. However, the model’s massive computational demands and large size necessitate expensive hardware, thus incurring significant costs in terms of training and inference. Moreover, recent T2I works [5; 6; 39; 1] do not release their training datasets, making it challenging for the open-source community to reproduce their performance due to internal or commercial restrictions.
@@
In this work, our goal is to build an efficient T2I model by exploring how far we can push the performance using only open-source data and models alone. To this end, we first design two compressed U-Nets, KOALA-1B and KOALA-700M, using not only block removal but also layer-wise removal to reduce the model size of the SDXL U-Net by up to 54% and 69% (vs. BK’s method: 33%) in Tab. 2. Then, we explore how to enhance the generative capability of the compact U-Net based on three empirical findings; (1) self-attention based knowledge distillation: we investigate how to effectively distill SDXL as a teacher model and find essential factors for feature-level KD. Specifically, we found that self-attention features are the most crucial for distillation since self-attention-based KD allows models to learn more discriminative representations between objects or attributes. (2) Data: When performing KD-training, high-resolution images and longer captions are more critical, even with fewer samples, than a larger number of low-resolution images with short captions. (3) Teacher model: The teacher model with higher performance improves the capability of the student model, and step-distilled teachers allow the student model to reduce the denoising steps, which results in further speed-up.
@@ 3 Three lessons for building an efficient text-to-image model
In this section, we introduce three empirical lessons to realize an efficient text-to-image synthesis; first, we design a lightweight U-Net architecture and perform a comprehensive analysis with the proposed efficient U-Net to find knowledge-distillation (KD) strategies in Sec. 3.1.2. Secondly, we investigate the training data characteristics that affect image generation quality in Sec. 3.2. Finally, we explore how different teacher models influence the student model in Sec. 3.3. For these empirical studies, we adopt two evaluation metrics, Human Preference Score (HPSv2) [54] and CompBench [17], instead of FID. Recently, several works [2; 54; 31] have claimed that FID [13] is not closely correlated with visual fidelity. HPSv2 [54] is for a visual aesthetics metric, which allows us to evaluate visual quality in terms of more specific types. As an image-text alignment metric, Compbench [17] is a more comprehensive benchmark for evaluating the compositional text-to-image generation capability than the single CLIP score [12]. We report average scores for HPS and Compbench, respectively.
@@ 3.1.2 Exploring Knowledge Distillation for SDXL
Self-attention-based knowledge distillation transfers discriminative image representation. With the increasing complexity of U-Net and its stage, relying solely on the last feature (LF) as in BK [21] may not be sufficient to mimic the intricate behavior of the teacher U-Net. Thus, we revisit which features provide the richest guidance for effective knowledge distillation. We focus on key intermediate features from each stage: outputs from the self-attention (SA), cross-attention (CA), and feedforward net (FFN) in the transformer block, as well as outputs from convolutional residual block (Res) and LF.  Tab. 3 summarizes the experimental results. While all types of features help obtain higher performance over the naïve baseline with only the task loss, distilling self-attention features achieves the most performance gain. Considering the prior studies [22; 46; 49] which suggest that SA plays a vital role in capturing semantic affinities and the overall structure of images, the results emphasize that such information is crucial for the distillation process.
@@
Self-attention at the decoder has a larger impact on the quality of generated images. We further explore the role and significance of each self-attention stage. To this end, we first visualize the self-attention map in  Fig. 3-(c). The self-attention maps initially capture general contextual information (e.g., DW-2&DW-3) and gradually focus on localized semantics (e.g., MID). In the decoder, self-attentions increasingly correlate with higher-level semantics (e.g., object) to accurately model appearances and structures. Notably, in this stage, the SA-based model attends corresponding object regions (given the query patch, red box) more discriminatively than the LF-based model, which results in improved compositional image generation performance.
@@ 4.1 Implementation details
Training. According to the recipe about the Teacher model in Sec. 3.3, we use SDXL-Turbo [39] and SDXL-Lightning [23] as teacher models, building two types of KOALA models, KOALA-Turbo and KOALA-Lightning. Since SDXL-Turbo and -Lightning are based on SDXL-Base model [31], we use the same two text encoders, OpenCLIP ViT-bigG [18] and CLIP ViT-L [33] and only replace the original U-Net with our efficient KOALA U-Net. Our U-Nets are initialized with the teacher’s U-Net weights at the exact block location. Using our self-attention-based KD method in Sec. 3.1.2, we train our KOALA models on the LAION-POP dataset using four NVIDIA A100 (80GB) GPUs with $512^{2}$ and $1024^{2}$ resolutions for KOALA-Turbo and KOALA-Lightning, respectively. Inference. We use Euler discrete scheduler [20] as the same sampler in SDXL [31]. All KOALA models generate images with 10 denoising steps, FP16, and cfg-sale [15] of 3.5. Please see further details in App. A.
@@ 5 Limitations
Figure 8: Failure cases of KOALA-700M
While our KOALA models generate images with decent aesthetic quality, such as photo-realistic or 3D-art renderings, they still show limitations in synthesizing legible texts in the generated image as shown in Fig. 8 (Left). Also, our models have difficulty in generating complex prompts with multiple attributed or object relationships, as shown in Fig. 8 (Right). Additionally, since SDXL is the de facto T2I model, we have tried to compress the SDXL U-Net by addressing its bottlenecks. However, this approach is somewhat specific to the SDXL U-Net and heuristic. This limitation arises because the SDXL U-Net has a complex and heterogeneous architecture, comprising both convolutional and transformer blocks, which hinders the formulation of a more general compression principle.
@@ 6 Conclusion
We used SDXL-Turbo [39] and SDXL-Lightning [23] as teacher models for feature guidance without applying their step-distillation training. For future work, since their methods and our KD approach are orthogonal, applying step-distillation to our KOALA backbone could yield synergistic effects, leading to significant speed improvements. Additionally, since we have identified the potential of applying our self-attention-based KD to Diffusion Transformer (DiT) models [30; 5; 6] in Tab. 8 due to their architectural simplicity compared to the U-Net in SDXL, we plan to further explore more general model compression methods for DiT, as in the language model literature [9; 28], and KD techniques based on our self-attention distillation.
@@ 7 Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training).
@@ A.1 Data
Synthesized captions, synCap. As some work [56; 1; 5] show that making caption data richer in information improves its generation quality, refining (cleaning) the training caption data and synthesizing more detailed captions corresponding to each paired image by a large multimodal model, e.g., LLaVA [25; 24] would improve the image-text alignment capability. To do this, we utilize LLaVA-1.5 [24] to synthesize captions corresponding to images in LAION-Aesthetics V2 6+. As shown in Fig. 9, when we input the instruction to LLaVA to describe the details of the input image, we can get highly augmented captions. It is worth noting that contrary to Pixart-$\alpha$ [5], which replaces original captions with synthesized ones, our approach appends augmented captions to the original ones, leveraging the contextual richness of existing proper nouns (e.g., New York).
@@ A.2 Training
Common training protocol. First, we describe a common training protocol for all experiments in our work. We base our framework on the officially released SDXL-Base-1.0333https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 and Diffusers library [51; 52]. We mainly replace computationally burdened SDXL’s U-Net with our efficient U-Net. We keep the same two text encoders, OpenCLIP ViT-bigG [18] and CLIP ViT-L [33], used in SDXL. For VAE, we use sdxl-vae-fp16-fix [4], which enables us to use FP16 precision for VAE computation. We initialize the weights of our U-Net with the teacher’s U-Net weights at the same block location. We freeze the text encoders, VAE, and the teacher U-Net of SDXL and only fine-tune our U-Net. When training, we use a discrete-time diffusion schedule [16], size- and crop-conditioning as in SDXL [31], AdamW optimizer [26], a batch size of 128, a constant learning rate of $10^{-5}$, and FP16 precision.
For the ablation study on the knowledge distillation strategies in Tabs. 3, 3 and 10, following the common training protocol except for batch size and training iteration, for fast verification, we train our KOALA models for 30k iterations with a batch size of 32 and $1024\times 1024$ resolution on LAION-Aesthetics V2 6+ [43] dataset using one NVIDIA A100 (80GB) GPU.

### v3

@@ 1 Introduction
Since Stable Diffusion XL [31] (SDXL) has become the de facto standard model for text-to-image (T2I) synthesis due to its ability to generate high-resolution images and its open-source nature, many models for specific downstream tasks [7; 47; 3; 58; 56] now leverage SDXL as their backbone. However, the model’s massive computational demands and large size necessitate expensive hardware, thus incurring significant costs in terms of training and inference. Moreover, recent T2I works [5; 6; 39; 1] do not release their training datasets, making it challenging for the open-source community to reproduce their performance due to internal or commercial restrictions.
@@
In this work, our goal is to build an efficient T2I model based on knowledge distillation in the latent diffusion model (KOALA) by exploring how far we can push the performance using only open-source data and models alone. To this end, we first design two compressed U-Nets, KOALA-1B and KOALA-700M, using not only block removal but also layer-wise removal to reduce the model size of the SDXL U-Net by up to 54% and 69% (vs. BK’s method: 33%) in Tab. 2. Then, we explore how to enhance the generative capability of the compact U-Net based on three empirical findings; (1) self-attention based knowledge distillation: we investigate how to effectively distill SDXL as a teacher model and find essential factors for feature-level KD. Specifically, we found that self-attention features are the most crucial for distillation since self-attention-based KD allows models to learn more discriminative representations between objects or attributes. (2) Data: When performing KD-training, high-resolution images and longer captions are more critical, even with fewer samples, than a larger number of low-resolution images with short captions. (3) Teacher model: The teacher model with higher performance improves the capability of the student model, and step-distilled teachers allow the student model to reduce the denoising steps, which results in further speed-up.
@@ 3 Three lessons for building an efficient text-to-image model
In this section, we introduce three empirical lessons to realize an efficient text-to-image synthesis; first, we design a lightweight U-Net architecture and perform a comprehensive analysis with the proposed efficient U-Net to find knowledge-distillation (KD) strategies in Sec. 3.1.2. Secondly, we investigate the training data characteristics that affect image generation quality in Sec. 3.2. Finally, we explore how different teacher models influence the student model in Sec. 3.3. For these empirical studies, we adopt two evaluation metrics, Human Preference Score (HPSv2) [55] and CompBench [17], instead of FID. Recently, several works [2; 55; 31] have claimed that FID [13] is not closely correlated with visual fidelity. HPSv2 [55] is for a visual aesthetics metric, which allows us to evaluate visual quality in terms of more specific types. As an image-text alignment metric, Compbench [17] is a more comprehensive benchmark for evaluating the compositional text-to-image generation capability than the single CLIP score [12]. We report average scores for HPS and Compbench, respectively.
@@ 3.1.2 Exploring Knowledge Distillation for SDXL
Self-attention-based knowledge distillation transfers discriminative image representation.
With the increasing complexity of U-Net and its stage, relying solely on the last feature (LF) as in BK [21] may not be sufficient to mimic the intricate behavior of the teacher U-Net. Thus, we revisit which features provide the richest guidance for effective knowledge distillation. We focus on key intermediate features from each stage: outputs from the self-attention (SA), cross-attention (CA), and feedforward net (FFN) in the transformer block, as well as outputs from convolutional residual block (Res) and LF.  Tab. 3 summarizes the experimental results. While all types of features help obtain higher performance over the naïve baseline with only the task loss, distilling self-attention features achieves the most performance gain. Considering the prior studies [22; 46; 49] which suggest that SA plays a vital role in capturing semantic affinities and the overall structure of images, the results emphasize that such information is crucial for the distillation process.
@@
Self-attention at the decoder has a larger impact on the quality of generated images.
We further explore the role and significance of each self-attention stage. To this end, we first visualize the self-attention map in  Fig. 3-(c). The self-attention maps initially capture general contextual information (e.g., DW-2&DW-3) and gradually focus on localized semantics (e.g., MID). In the decoder, self-attentions increasingly correlate with higher-level semantics (e.g., object) to accurately model appearances and structures. Notably, in this stage, the SA-based model attends corresponding object regions (given the query patch, red box) more discriminatively than the LF-based model, which results in improved compositional image generation performance.
@@ 4.1 Implementation details
Training. According to the recipe about the Teacher model in Sec. 3.3, we use SDXL-Turbo [39] and SDXL-Lightning [23] as teacher models, building two types of KOALA models, KOALA-Turbo and KOALA-Lightning. Since SDXL-Turbo and -Lightning are based on SDXL-Base model [31], we use the same two text encoders, OpenCLIP ViT-bigG [18] and CLIP ViT-L [33] and only replace the original U-Net with our efficient KOALA U-Net. Our U-Nets are initialized with the teacher’s U-Net weights at the exact block location. Using our self-attention-based KD method in Sec. 3.1.2, we train our KOALA models on the LAION-POP dataset using four NVIDIA A100 (80GB) GPUs with $512^{2}$ and $1024^{2}$ resolutions for KOALA-Turbo and KOALA-Lightning, respectively. Inference. We use Euler discrete scheduler [20] as the same sampler in SDXL [31]. All KOALA models generate images with 10 denoising steps, FP16, and cfg-sale [15] of 3.5. Please see further details of training and inference in Sec. A.2 and Sec. A.3, respectively.
@@ 4.3 Discussion
Synergy Effect with Step-Distillation Method, PCM [53]. Since step-distillation methods [23; 39; 53] and our KD approach are orthogonal, applying our KOALA backbones to the step-distillation methods could yield synergistic effects, leading to further speed improvements. To verify the synergy effect between the step-distillation method (e.g., PCM [53]) and our KOALA backbones, we conduct step-distillation training using PCM with our KOALA backbones, and the results are presented in Tab. 9. Thanks to their efficiency, our KOALA backbones allow PCM [^2] PCM is the only work that officially provides step-distillation training code, [^3] https://github.com/G-U-N/Phased-Consistency-Model to achieve additional speed-up with only a slight performance drop compared to using the SDXL-Base backbone. Furthermore, we provide qualitative comparisons between PCM-KOALA models and PCM-SDXL-Base in Fig. 20, demonstrating that the generated images achieve visual quality comparable to those of SDXL-Base.
@@
Table 9: Synergy effect with a step-distilled method, PCM [53]. We conduct step-distillation training using PCM with our KOALA backbones and compare with PCM-SDXL-Base.
Method Teacher Student #Step Param. (B) Memory (GB) Latency HPS CompBench PCM [53] SDXL-Base SDXL-Base 2 2.56 12.1 0.345 29.99 0.4169 PCM [53] SDXL-Base KOALA-700M 2 0.78 8.2 0.222 28.78 0.3930 PCM [53] SDXL-Base KOALA-1B 2 1.16 9.0 0.235 29.04 0.4055
@@ 5 Limitations
While our KOALA models generate images with decent aesthetic quality, such as photo-realistic or 3D-art renderings, they still show limitations in synthesizing legible texts in the generated image as shown in Fig. 8 (Left). Also, our models have difficulty in generating complex prompts with multiple attributed or object relationships, as shown in Fig. 8 (Right). Additionally, since SDXL is the de facto T2I model, we have tried to compress the SDXL U-Net by addressing its bottlenecks. However, this approach is somewhat specific to the SDXL U-Net and heuristic. This limitation arises because the SDXL U-Net has a complex and heterogeneous architecture, comprising both convolutional and transformer blocks, which hinders the formulation of a more general compression principle. More detailed investigations and examples are described in App. D.
@@ 6 Conclusion
Additionally, since we have identified the potential of applying our self-attention-based KD to Diffusion Transformer (DiT) models [30; 5; 6] in Tab. 8 due to their architectural simplicity compared to the U-Net in SDXL, we plan to further explore more general model compression methods for DiT, as in the language model literature [9; 28], and KD techniques based on our self-attention distillation.
@@ 7 Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training, 45%), (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration, 45%) and (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST), 10%).
@@ A.1 Data
Synthesized captions, synCap. As some work [57; 1; 5] show that making caption data richer in information improves its generation quality, refining (cleaning) the training caption data and synthesizing more detailed captions corresponding to each paired image by a large multimodal model, e.g., LLaVA [25; 24] would improve the image-text alignment capability. To do this, we utilize LLaVA-1.5 [24] to synthesize captions corresponding to images in LAION-Aesthetics V2 6+. As shown in Fig. 9, when we input the instruction to LLaVA to describe the details of the input image, we can get highly augmented captions. It is worth noting that contrary to Pixart-$\alpha$ [5], which replaces original captions with synthesized ones, our approach appends augmented captions to the original ones, leveraging the contextual richness of existing proper nouns (e.g., New York).
@@ A.2 Training
Common training protocol. First, we describe a common training protocol for all experiments in our work. We base our framework on the officially released SDXL-Base-1.0555https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 and Diffusers library [51; 52]. We mainly replace computationally burdened SDXL’s U-Net with our efficient U-Net. We keep the same two text encoders, OpenCLIP ViT-bigG [18] and CLIP ViT-L [33], used in SDXL. For VAE, we use sdxl-vae-fp16-fix [4], which enables us to use FP16 precision for VAE computation. We initialize the weights of our U-Net with the teacher’s U-Net weights at the same block location. We freeze the text encoders, VAE, and the teacher U-Net of SDXL and only fine-tune our U-Net. When training, we use a discrete-time diffusion schedule [16], size- and crop-conditioning as in SDXL [31], AdamW optimizer [26], a batch size of 128, a constant learning rate of $10^{-5}$, and FP16 precision.
For the ablation study on the knowledge distillation strategies in Tabs. 3, 3 and 11, following the common training protocol except for batch size and training iteration, for fast verification, we train our KOALA models for 30k iterations with a batch size of 32 and $1024\times 1024$ resolution on LAION-Aesthetics V2 6+ [43] dataset using one NVIDIA A100 (80GB) GPU.
@@ Appendix E Societal Impacts
Figure 20: Qualitative comparison between PCM-SDXL and our PCM-KOALA models with $1024^{2}$.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We made sure that the claims made in the abstract and introduction accurately reflect our contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of our method in Sec. 5, including failure cases generated by our KOALA and our simple and specific compression method.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed instructions of how to reproduce all of the experiments, hyperparameters, and the models and datasets used, all of which are open-source.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We already provided where we can download the training datasets and evaluation benchmarks. However, we cannot provide the training code and checkpoint weights due to the internal policy of our company. After getting permission from our company, we will release the trained weights in Huggingface to generate images.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We include details for dataset preparation and experimental settings in App. A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: In our study, we compared various models using quantitative evaluations, specifically focusing on HPSv2 and Compbench. The evaluation of HPSv2 required generating 2,400 images, while Compbench required 24,000 images, resulting in a significant inference cost that made repeated inferences challenging. Despite this, we believe that our performance evaluations are reliable as the performance deviations are expected to be minimal based on the sufficient number of test images used in our evaluations. Furthermore, we did not cherry-pick any images during the quantitative evaluation due to the sheer volume of images generated and evaluated. This extensive evaluation approach ensures that the reported performance metrics are representative and robust.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the relevant details in App. A and Sec. 4.4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have thoroughly reviewed the ethics guidelines and we have made sure to preserve anonymity for the review process.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We already addressed the broader impacts in App. E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: As a text-to-image generation model, our KOALA, addresses the potential risks of generating Not Safe For Work (NSFW) content, which includes harmful, violent, or adult imagery. To mitigate these risks, KOALA integrates NSFW content detection capabilities provided by Huggingface and the transformers library. By calculating the NSFW score for each generated image and filtering out those that exceed a predefined threshold, our model effectively prevents the creation and dissemination of inappropriate content. This approach ensures the ethical and responsible use of AI technology, aligning our contributions with societal values and ethical guidelines.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: we have clearly cited the teacher models used for training KOALA: SDXL-Base, SDXL-Turbo, and SDXL-Lightning in the main paper and Appendix.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We didn’t provide new assets except for our main paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We have not used crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We have not used crowdsourcing or human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.